# SVFAP: Self-supervised Video Facial Affect Perceiver

### v1

@@ Abstract
Video-based facial affect analysis has recently attracted increasing attention owing to its critical role in human-computer interaction. Previous studies mainly focus on developing various deep learning architectures and training them in a fully supervised manner. Although significant progress has been achieved by these supervised methods, the longstanding lack of large-scale high-quality labeled data severely hinders their further improvements. Motivated by the recent success of self-supervised learning in computer vision, this paper introduces a self-supervised approach, termed Self-supervised Video Facial Affect Perceiver (SVFAP), to address the dilemma faced by supervised methods. Specifically, SVFAP leverages masked facial video autoencoding to perform self-supervised pre-training on massive unlabeled facial videos. Considering that large spatiotemporal redundancy exists in facial videos, we propose a novel temporal pyramid and spatial bottleneck Transformer as the encoder of SVFAP, which not only enjoys low computational cost but also achieves excellent performance. To verify the effectiveness of our method, we conduct experiments on nine datasets spanning three downstream tasks, including dynamic facial expression recognition, dimensional emotion recognition, and personality recognition. Comprehensive results demonstrate that SVFAP can learn powerful affect-related representations via large-scale self-supervised pre-training and it significantly outperforms previous state-of-the-art methods on all datasets. Codes will be available at https://github.com/sunlicai/SVFAP.
@@ 1 Introduction
Video-based facial affect analysis, which aims to automatically detect and understand human affective states from facial videos, has recently gained considerable attention due to its great potential in developing natural and harmonious human-computer interaction systems [1, 2, 3]. Early attempts for this task focus on designing advanced handcrafted features and machine learning algorithms on small lab-controlled datasets. With the advent of deep learning and larger labeled datasets, the research paradigm has changed to train supervised deep neural networks in an end-to-end manner. Researchers have developed a variety of deep architectures to improve model performance, including convolutional neural networks (CNN) [4, 5], recurrent neural networks (RNN) [6, 7], Transformers [8, 9], and their combinations [10, 5, 11, 12, 13].
@@
Despite its great success in many deep learning fields, self-supervised learning has rarely been explored in video-based affective computing. To this end, this paper presents a self-supervised learning method, named Self-supervised Video Facial Affect Perceiver (SVFAP), to unleash the power of large-scale self-supervised learning for video-based facial affect analysis. As shown in Fig. 1, SVFAP involves two-stage training, i.e., self-supervised pre-training and downstream fine-tuning. Its whole pipeline is conceptually simple and generally inherits those of MAE [22] and its video versions [23, 24] considering their big success in computer vision. Concretely, during the pre-training stage, SVFAP utilizes masked facial video autoencoding to learn useful and transferable spatiotemporal representations from a large amount of unlabeled facial video data. The architecture adopts an asymmetric encoder-decoder design [22] to enable efficient pre-training, in which a high-capacity encoder only processes limited visible input (since the masking ratio is very high) and a lightweight decoder operates all input and reconstructs the masked part. During fine-tuning, it discards the decoder and fine-tunes the pre-trained high-capacity encoder on downstream datasets. Note that, since the vanilla Vision Transformer (ViT) [25] is typically employed as the encoder in MAE and its video versions, the computational cost is very expensive during downstream fine-tuning (especially for videos) despite the architecture efficiency in pre-training. Considering that large redundancy (e.g., facial symmetry and temporal correlation) exists in 3D facial video data, we thus propose a novel Temporal Pyramid and Spatial Bottleneck Transformer (TPSBT) as the high-capacity encoder of SVFAP to achieve both efficient pre-training and fine-tuning. As the name suggests, TPSBT (Fig. 2) utilizes spatial bottleneck mechanism and temporal pyramid learning to minimize redundant information in spatial and temporal dimensions, which not only reduces the computational cost greatly (about 43% FLOPs reduction during fine-tuning) but also leads to superior performance.
@@
To verify the effectiveness of SVFAP, we conduct experiments on nine datasets from three video-based facial affect analysis tasks, including six datasets for dynamic facial expression recognition, two datasets for dimensional emotion recognition, and one dataset for personality recognition. Comprehensive experimental results demonstrate that our SVFAP can learn powerful affect-related representations from large-scale unlabeled facial video data via self-supervised pre-training and significantly outperforms previous state-of-the-art methods on all downstream datasets. For instance, on three in-the-wild dynamic facial expression recognition datasets, our best model surpasses the previous best by 5.72% UAR and 5.03% WAR on on DFEW [5], 4.38% UAR and 3.85% WAR on FERV39k [26], and 7.91% UAR and 6.10% WAR on MAFW [27]. To sum up, the main contributions of this paper are three-fold:
@@
  • We propose a novel TPSBT model as the encoder of SVFAP to eliminate large spatiotemporal redundancy in facial videos, which not only enjoys lower computational cost but also has superior performance when compared with vanilla ViT.
@@ 2.1 Video-based Facial Affect Analysis
Most studies on video-based facial affect analysis fall into the supervised learning paradigm. They mainly concentrate on developing more advanced deep learning architectures to extract discriminative spatiotemporal representations from raw facial videos. Generally, there are two lines of research. The first line of research treats the spatial and temporal dimension of 3D video data in an independent manner [10, 5, 11, 27, 26, 28]. Typically, 2D CNN (e.g., VGGNet [29] and ResNet [30]) is first used to extract facial features from each static frame and then RNN (e.g., LSTM [31] and GRU [32]) runs over them to integrate the dynamic temporal information across all frames. Recently, inspired by the great success of Transformer [33] in natural language processing and computer vision, there are also several studies that utilize the global dependency modeling ability of Transformer to enhance spatial and temporal feature extraction of traditional CNN and RNN [12, 9, 13, 8, 34]. Another line of research tries to simultaneously encode spatial appearance features and dynamic motion information by extending the 2D convolution kernel to the 3D convolution kernel along the temporal axis. With the help of 3D kernels, 3D CNN (e.g., C3D [35], R(2+1)D [36], P3D [37], and 3D ResNet [38]) is expected to extract discriminative spatiotemporal representations from raw videos.
@@
Figure 1: An overview of the proposed method (i.e, SVFAP). It consists of two stages, including self-supervised pre-training (a) and downstream fine-tuning (b). During pre-training, SVFAP utilizes masked facial video autoencoding as the training objective. Following previous studies [22, 23, 24], it adopts an asymmetric encoder-decoder architecture and a high masking ratio (e.g., 90%) to enable fast pre-training on large-scale unlabeled facial video data. After pre-training, the lightweight decoder is discarded and the high-capacity encoder is fine-tuned in downstream tasks.
@@ 2.2 Self-supervised Learning
Self-supervised learning aims to solve the data-hungry issue in supervised learning by exploiting massive unlabeled data. It can be roughly divided into two categories: discriminative and generative [20]. The discriminative method generally follows the supervised counterpart by designing a discriminative loss. Early studies focused on developing various geometry-based pretext tasks, such as predicting image rotation [39] and sorting shuffled video frames [40]. In recent years, the trend has shifted from handcrafted tasks to contrastive learning methods, e.g., MoCo [41], and SwAV [42]. Until the recent reviving success of generative methods, contrastive learning methods are the dominant self-supervised pre-training framework in computer vision. The generative method typically involves an autoencoding process, in which an encoder maps the input into a latent representation and a decoder reconstructs the original input from encoded representation. Early generative work can be dated back to denoising autoencoder [43], i.e., reconstructing the clean input from a partially corrupted one. Recently, motivated by the unprecedented success in natural language processing (e.g., BERT [21] and GPT [44]), generative methods also achieve impressive results in computer vision [45, 22]. One of the most representative methods is masked autoencoder (MAE) [22]. MAE and its video versions [23, 24] utilize an asymmetric encoder-decoder architecture to efficiently pre-train vanilla ViT and outperforms its supervised counterpart and state-of-the-art contrastive learning methods by large margins in many downstream tasks.
@@
Despite the great success of self-supervised learning in many deep learning fields, it has rarely been explored in video-based facial affect analysis. Roy et al. [46] present a spatiotemporal contrastive learning method pre-trained on a small lab-controlled labeled dataset. Unlike it, the proposed SVFAP is built upon more advanced generative methods (i.e., MAE and and its video versions) and leverages large-scale in-the-wild unlabeled facial videos for self-supervised pre-training. There are also several self-supervised studies in relevant research field (i.e., facial action unit detection). For instance, FAb-Net [47] exploits relative facial movements between adjacent frames as free supervisory signal to perform self-supervised pre-training. TCAE [48] improves FAb-Net by disentangling the head pose related movements and facial action related ones. FaceCycle [49] further introduces facial motion and identity cycle-consistency to promote facial representation learning. Lu et al. [50] proposes a triplet-based frame ranking method for temporal consistency modeling. Recently, there are also a few studies focusing on general self-supervised facial representation learning. Bulat et al. [51] pre-train SwAV [42] on large-scale face recognition datasets and find that it achieves significant improvements over previous supervised methods on five face analysis tasks. FaRL [52] utilizes both low-level masked facial image modeling and high-level face-text contrastive learning for self-supervised pre-training and outperforms state-of-the-art methods on many face analysis tasks. Although achieving promising results, these methods use 2D models for self-supervised pre-training, thus could not capture rich spatiotemporal information in 3D facial videos.
@@ 2.3 Vision Transformer
Originated from natural language processing, Transformer [33] based architectures have recently revolutionized various computer vision tasks by means of its strong long-range dependency modeling ability [53]. Among them, the pioneering work of ViT [25] directly applies the standard Transformer to a sequence of image patches and performs very well on image classification tasks, challenging the dominant paradigm of CNN in computer vision. Since ViT relies on large amounts of labeled data (i.e., JFT-300M) to achieve successful supervised pre-training, DeiT [54] introduces several training strategies to allow it can be trained on much smaller ImageNet-1K dataset. After that, numerous ViT variants have emerged by incorporating more or less vision-friendly priors (e.g., local self-attention and hierarchical design) to reduce the quadratic scaling cost of vanilla ViT and improve model performance in vision tasks [55, 56, 57, 58, 59]. For instance, in the video domain, TimeSformer [56] applies factorised temporal and spatial attention in the standard Transformer block to achieve spatiotemporal representation learning. MViT [57] introduces multi-scale feature hierarchies to the vanilla ViT. Video Swin Transformer[59] utilizes 3D shifted window attention to inject spatiotemporal locality inductive bias into video Transformers. Although these variants usually performs better than vanilla ViT in the supervised setting, they are typically not suitable for masked autoencoding in self-supervised pre-training. This is because they could not drop masked tokens due to the incorporation of local operations (e.g., window-based attention and patch merging) in their architectures. Thus, they can not enjoy the efficiency of vanilla ViT during masked autoencoding. This is why MAE and its video versions all employ ViT as the encoder. Nevertheless, ViT still suffers from expensive computational cost during downstream fine-tuning. To this end, in this paper, we introduce temporal pyramid learning and spatial bottleneck mechanism to ViT to enable both efficient pre-training and fine-tuning.
@@ 3.1 Self-supervised Pre-training
The idea of masked facial video autoencoding is simple, i.e., reconstructing the original facial video given partially observed input. Concretely, SVFAP consists of an encoder to project the original video to a high-level spatiotemporal representation in the latent space, and then a decoder to reconstruct the original input from the encoded representation. Moreover, similar to MAE [22] and its video versions [23, 24], SVFAP adopts an asymmetric encoder-decoder architecture, where a high-capacity encoder only accepts limited visible tokens (e.g., 10%) as input and a lightweight decoder operates on all tokens to perform the video reconstruction. With this special design, the computational cost could be largely reduced to allow fast pre-training.
@@ 3.1.3 Encoder
Typically, the vanilla ViT [25] is employed as the encoder in masked autoencoding [22, 23, 24]. Although it enjoys much efficiency during pre-training by discarding many masked tokens, its computational cost is still very expensive during downstream fine-tuning as it must take as input all embedded tokens. Meanwhile, we notice that large redundancy exists in 3D facial videos, such as facial symmetry and temporal correlation. Motivated by this, we propose the Temporal Pyramid and Spatial Bottleneck Transformer (TPSBT) as the encoder of our SVFAP to enable both fast pre-training and fine-tuning. As shown in Fig. 2, our TPSBT has three stages. The first stage is composed of several standard Transformer blocks, while in the last two stages, they are replaced by spatial bottleneck Transformer blocks to reduce spatial redundancy. Moreover, benefiting from tube masking, we could further utilize temporal pyramid learning to reduce redundancy in the temporal dimension. For convenience, in the following part, we denote as $\mathbf{X}\in\mathbb{R}^{N\times C}$ the reshaped input token embeddings after tube masking, where $N=\frac{T}{2}\cdot S$ is the total number of visible tokens and $S=\frac{H}{16}\cdot\frac{W}{16}\cdot(1-\rho$) denotes the spatial token number. And $\mathbf{X}_{i}\in\mathbb{R}^{N_{i}\times C}$ represents the output of stage $i$ ($N_{i}$ is the number of tokens in each stage, $i\in\{1,2,3\}$).
Standard Transformer. The standard Transformer in stage 1 is used to retain the global spatiotemporal representation learning ability of vanilla ViT. It consists of a sequence of $M_{1}$ blocks, each of which is composed of alternating layers of Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN), with Layer Normalization (LN) applied before each layer and residual connections after each layer. Formally, we define a standard Transformer block as follows:
@@
where $\mathbf{Q}_{h}=\mathbf{X}\mathbf{W}^{Q}_{h}$ is the query, $\mathbf{K}_{h}=\mathbf{X}\mathbf{W}^{K}_{h}$ is the key, $\mathbf{V}_{h}=\mathbf{X}\mathbf{W}^{V}_{h}$ is the value, $\mathbf{W}^{*}_{h}\in\mathbb{R}^{C\times d_{h}}$ (*$*\in\{Q,K,V\}$), $\mathbf{W}^{O}\in\mathbb{R}^{C\times C}$, $H$ is the number of heads, $d_{h}=C/H$ is the feature dimension of head $h$, and the superscript ${}^{\top}$ is matrix transpose. Finally, FFN in Eq. (1) consists of two linear layers with a GELU [60] non-linearity in between, i.e.,
@@
Subsequently, we utilizes $M_{i}-1$ SBT blocks to operate on the summarized bottleneck tokens to achieve efficient global spatiotemporal interactions. As shown in Fig. 2, an SBT block mainly consists of a Multi-Head Cross-Attention (MHCA) layer [61], an MHSA layer, and an FFN layer. Formally, it can be defined as follows:
@@
where the superscript $j\in\{1,...,M_{i}-1\}$ is the block index, $\mathbf{B}^{(0)}_{i}=\mathbf{B}_{i}$. The MHCA layer iteratively distills necessary information from original input back into the bottleneck embeddings to avoid key information lost during spatial compression. It is a variant of MHSA, whose query is the bottleneck embeddings while the key and value come from the original input to allow information flow from the latter to the former, i.e.,
@@
Compared with a standard Transformer block, the computational complexity of an SBT block is reduced from the quadratic $O(S^{2}T_{i}^{2})$ to $O(G(G+S)T_{i}^{2})$ thanks to the introduction of spatial bottleneck tokens. Considering that $G$ is a small constant, an SBT block thus enjoys linear computational complexity with respect to the spatial size $S$. Note that, empirically, we have $G\approx S$ during pre-training and $G\ll S$ during fine-tuning (Sec. 4.1). Therefore, the computational cost of pre-training remains approximately unchanged (actually, slightly decreases) and large computations can be reduced during downstream fine-tuning, allowing both efficient pre-training and fine-tuning (Sec. 4.2.2).

### v2

@@ Abstract
Video-based facial affect analysis has recently attracted increasing attention owing to its critical role in human-computer interaction. Previous studies mainly focus on developing various deep learning architectures and training them in a fully supervised manner. Although significant progress has been achieved by these supervised methods, the longstanding lack of large-scale high-quality labeled data severely hinders their further improvements. Motivated by the recent success of self-supervised learning in computer vision, this paper introduces a self-supervised approach, termed Self-supervised Video Facial Affect Perceiver (SVFAP), to address the dilemma faced by supervised methods. Specifically, SVFAP leverages masked facial video autoencoding to perform self-supervised pre-training on massive unlabeled facial videos. Considering that large spatiotemporal redundancy exists in facial videos, we propose a novel temporal pyramid and spatial bottleneck Transformer as the encoder of SVFAP, which not only largely reduces computational costs but also achieves excellent performance. To verify the effectiveness of our method, we conduct experiments on nine datasets spanning three downstream tasks, including dynamic facial expression recognition, dimensional emotion recognition, and personality recognition. Comprehensive results demonstrate that SVFAP can learn powerful affect-related representations via large-scale self-supervised pre-training and it significantly outperforms previous state-of-the-art methods on all datasets. Code is available at https://github.com/sunlicai/SVFAP.
@@ 1 Introduction
Video-based facial affect analysis, which aims to automatically detect and understand human affective states from facial videos, has recently gained considerable attention due to its great potential in developing natural and harmonious human-computer interaction systems [1, 2, 3, 4]. Early attempts for this task focus on designing advanced handcrafted features and machine learning algorithms on small lab-controlled datasets. With the advent of deep learning and larger labeled datasets, the research paradigm has changed to train supervised deep neural networks in an end-to-end manner. Researchers have developed a variety of deep architectures to improve model performance, including convolutional neural networks (CNN) [5, 6], recurrent neural networks (RNN) [7, 8], Transformers [9, 10], and their combinations [11, 6, 12, 13, 14].
@@
Despite its great success in many deep learning fields, self-supervised learning has rarely been explored in video-based affective computing. To this end, this paper presents a self-supervised learning method, named Self-supervised Video Facial Affect Perceiver (SVFAP), to unleash the power of large-scale self-supervised learning for video-based facial affect analysis. As shown in Fig. 1, SVFAP involves two-stage training, i.e., self-supervised pre-training and downstream fine-tuning. Its whole pipeline is conceptually simple and generally inherits those of MAE [23] and its video versions [24, 25] considering their big success in computer vision. Concretely, during the pre-training stage, SVFAP utilizes masked facial video autoencoding to learn useful and transferable spatiotemporal representations from a large amount of unlabeled facial video data. The architecture adopts an asymmetric encoder-decoder design [23] to enable efficient pre-training, in which a high-capacity encoder only processes limited visible input (since the masking ratio is very high) and a lightweight decoder operates all input and reconstructs the masked part. During fine-tuning, it discards the decoder and fine-tunes the pre-trained high-capacity encoder on downstream datasets. Note that, since the vanilla Vision Transformer (ViT) [26] is typically employed as the encoder in MAE and its video versions, the computational costs are very expensive during downstream fine-tuning (especially for videos) despite the architecture efficiency in pre-training. Considering that large redundancy (e.g., facial symmetry and temporal correlation) exists in 3D facial video data, we thus propose a novel Temporal Pyramid and Spatial Bottleneck Transformer (TPSBT) as the high-capacity encoder of SVFAP to achieve both efficient pre-training and fine-tuning. As the name suggests, TPSBT (Fig. 2) utilizes spatial bottleneck mechanism and temporal pyramid learning to minimize redundant information in spatial and temporal dimensions, which not only reduces the computational costs greatly (about 43% FLOPs reduction during fine-tuning) but also leads to superior performance.
@@
To verify the effectiveness of SVFAP, we conduct experiments on nine datasets from three video-based facial affect analysis tasks, including six datasets for dynamic facial expression recognition, two datasets for dimensional emotion recognition, and one dataset for personality recognition. Comprehensive experimental results demonstrate that our SVFAP can learn powerful affect-related representations from large-scale unlabeled facial video data via self-supervised pre-training and significantly outperforms previous state-of-the-art methods on all downstream datasets. For instance, on three in-the-wild dynamic facial expression recognition datasets, our best model surpasses the previous best by 5.72% UAR and 5.02% WAR on DFEW [6], 4.38% UAR and 3.75% WAR on FERV39k [27], and 7.91% UAR and 6.10% WAR on MAFW [28]. To sum up, the main contributions of this paper are three-fold:
@@
  • We propose a novel TPSBT model as the encoder of SVFAP to eliminate large spatiotemporal redundancy in facial videos, which not only enjoys lower computational costs but also has superior performance when compared with the vanilla ViT.
@@ 2.1 Video-based Facial Affect Analysis
Most studies on video-based facial affect analysis fall into the supervised learning paradigm. They mainly concentrate on developing more advanced deep learning architectures to extract discriminative spatiotemporal representations from raw facial videos. Generally, there are two lines of research. The first line of research treats the spatial and temporal dimension of 3D video data in an independent manner [11, 6, 12, 28, 27, 29]. Typically, 2D CNN (e.g., VGGNet [30] and ResNet [31]) is first used to extract spatial features from each static frame and then RNN (e.g., LSTM [32] and GRU [33]) runs over them to integrate the dynamic temporal information across all frames. Recently, inspired by the great success of Transformer [34] in natural language processing and computer vision, there are also several studies that utilize the global dependency modeling ability of Transformer to enhance spatial and temporal feature extraction of traditional CNN and RNN [13, 10, 14, 9, 35]. Another line of research tries to simultaneously encode spatial appearance features and dynamic motion information by extending the 2D convolution kernel to the 3D convolution kernel along the temporal axis. With the help of 3D kernels, 3D CNN (e.g., C3D [36], R(2+1)D [37], P3D [38], and 3D ResNet [39]) is expected to extract discriminative spatiotemporal representations from raw videos.
@@ 2.2 Self-supervised Learning
Self-supervised learning aims to solve the data-hungry issue in supervised learning by exploiting massive unlabeled data. It can be roughly divided into two categories: discriminative and generative [21]. The discriminative method generally follows the supervised counterpart by designing a discriminative loss. Early studies focused on developing various geometry-based pretext tasks, such as predicting image rotation [40] and sorting shuffled video frames [41]. In recent years, the trend has shifted from handcrafted tasks to contrastive learning methods, e.g., MoCo [42], and SwAV [43]. Contrastive learning has been the dominant self-supervised pre-training framework in computer vision until the more recent reviving success of generative learning. The generative method typically involves an autoencoding process, in which an encoder maps the input into a latent representation and a decoder reconstructs the original input from the encoded representation. Early generative work can be dated back to denoising autoencoder [44], i.e., reconstructing the clean input from a partially corrupted one. Recently, motivated by the unprecedented success in natural language processing (e.g., BERT [22] and GPT [45]), generative methods have also achieved impressive results in computer vision [46, 23]. One of the most representative methods is masked autoencoder (MAE) [23]. MAE and its video versions [24, 25] utilize an asymmetric encoder-decoder architecture to efficiently pre-train vanilla ViT and outperform its supervised counterpart and state-of-the-art contrastive learning methods by large margins in many downstream tasks (e.g., object recognition/detection/segmentation, and action recognition/detection).
@@
Despite the great success of self-supervised learning in many deep learning fields, it has rarely been explored in video-based facial affect analysis. Roy et al. [47] present a spatiotemporal contrastive learning method pre-trained on a small lab-controlled labeled dataset. Unlike it, the proposed SVFAP is built upon more advanced generative methods (i.e., MAE and its video versions) and leverages large-scale in-the-wild unlabeled facial videos for self-supervised pre-training. There are also several self-supervised studies in the relevant research field (i.e., facial action unit detection). For instance, FAb-Net [48] exploits relative facial movements between adjacent frames as free supervisory signals to perform self-supervised pre-training. TCAE [49] improves FAb-Net by disentangling the head pose-related movements and facial action-related ones. FaceCycle [50] further introduces facial motion and identity cycle-consistency to promote facial representation learning. Lu et al. [51] propose a triplet-based frame ranking method for temporal consistency modeling. Recently, there are also a few studies focusing on general self-supervised facial representation learning. Bulat et al. [52] pre-train SwAV [43] on large-scale face recognition datasets and find that it achieves significant improvements over previous supervised methods on five face analysis tasks. FaRL [53] utilizes both low-level masked facial image modeling and high-level face-text contrastive learning for self-supervised pre-training and outperforms state-of-the-art methods on many face analysis tasks. Although achieving promising results, these methods use 2D models for self-supervised pre-training, thus could not capture rich spatiotemporal information in 3D facial videos.
@@ 2.3 Vision Transformer
Originated from natural language processing, Transformer [34] based architectures have recently revolutionized various computer vision tasks by means of its strong long-range dependency modeling ability [54]. Among them, the pioneering work of ViT [26] directly applies the standard Transformer to a sequence of image patches and performs very well on image classification tasks, challenging the dominant paradigm of CNN in computer vision. Since ViT relies on large amounts of labeled data (i.e., JFT-300M) to achieve successful supervised pre-training, DeiT [55] introduces several training strategies to allow it can be trained on a much smaller dataset (i.e., ImageNet-1K). After that, numerous ViT variants have emerged by incorporating more or less vision-friendly priors (e.g., local self-attention and hierarchical design) to reduce the quadratic scaling cost of vanilla ViT and improve model performance in vision tasks [56, 57, 58, 59, 60]. For instance, in the video domain, TimeSformer [57] applies factorized temporal and spatial attention in the standard Transformer block to achieve spatiotemporal representation learning. MViT [58] introduces multi-scale feature hierarchies to the vanilla ViT. Video Swin Transformer[60] utilizes 3D-shifted window attention to inject spatiotemporal locality inductive bias into video Transformers. Although these variants usually perform better than vanilla ViT in the supervised setting, they are typically not suitable for masked autoencoding in self-supervised pre-training. This is because they could not drop masked tokens due to the incorporation of local operations (e.g., window-based attention and patch merging) in their architectures. Thus, they can not enjoy the efficiency of vanilla ViT during masked autoencoding. This is why MAE and its video versions all employ ViT as the encoder. Nevertheless, ViT still suffers from expensive computational costs during downstream fine-tuning. To this end, in this paper, we introduce temporal pyramid learning and spatial bottleneck mechanism to ViT to enable both efficient pre-training and fine-tuning.
@@ 3 Method
Figure 1: An overview of the proposed method (i.e., SVFAP). It consists of two stages, including self-supervised pre-training (a) and downstream fine-tuning (b). During pre-training, SVFAP utilizes masked facial video autoencoding as the training objective. Following previous studies [23, 24, 25], it adopts an asymmetric encoder-decoder architecture and a high masking ratio (e.g., 90%) to enable fast pre-training on large-scale unlabeled facial video data. After pre-training, the lightweight decoder is discarded and the high-capacity encoder is fine-tuned in downstream tasks.
@@ 3.1 Self-supervised Pre-training
The idea of masked facial video autoencoding is simple, i.e., reconstructing the original facial video given partially observed input. Concretely, SVFAP consists of an encoder to project the original video to a high-level spatiotemporal representation in the latent space, and then a decoder to reconstruct the original input from the encoded representation. Moreover, similar to MAE [23] and its video versions [24, 25], SVFAP adopts an asymmetric encoder-decoder architecture, where a high-capacity encoder only accepts limited visible tokens (e.g., 10%) as input and a lightweight decoder operates on all tokens to perform the video reconstruction. With this special design, the computational costs could be largely reduced to allow fast pre-training.
@@ 3.1.3 Encoder
Typically, the vanilla ViT [26] is employed as the encoder in masked autoencoding [23, 24, 25]. Although it enjoys much efficiency during pre-training by discarding many masked tokens, its computational costs are still very expensive during downstream fine-tuning as it must take as input all embedded tokens. Meanwhile, we notice that large redundancy exists in 3D facial videos, such as facial symmetry and temporal correlation. Motivated by this, we propose the Temporal Pyramid and Spatial Bottleneck Transformer (TPSBT) as the encoder of our SVFAP to enable both fast pre-training and fine-tuning. As shown in Fig. 2, our TPSBT has three stages. The first stage is composed of several standard Transformer blocks, while in the last two stages, they are replaced by spatial bottleneck Transformer blocks to reduce spatial redundancy. Moreover, benefiting from tube masking, we could further utilize temporal pyramid learning to reduce redundancy in the temporal dimension. For convenience, in the following part, we denote as $\mathbf{X}\in\mathbb{R}^{N\times C}$ the reshaped input token embeddings after tube masking, where $N=\frac{T}{2}\cdot S$ is the total number of visible tokens and $S=\frac{H}{16}\cdot\frac{W}{16}\cdot(1-\rho$) denotes the spatial token number. And $\mathbf{X}_{i}\in\mathbb{R}^{N_{i}\times C}$ represents the output of stage $i$ ($N_{i}$ is the number of tokens in each stage, $i\in\{1,2,3\}$).
Standard Transformer. The standard Transformer in stage 1 is used to retain the global spatiotemporal representation learning ability of vanilla ViT. It consists of a sequence of $M_{1}$ Transformer blocks, each of which is composed of alternating layers of Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN), with Layer Normalization (LN) applied before each layer and residual connections after each layer. Formally, we define a standard Transformer block as follows:
@@
where $\mathbf{Q}_{h}=\mathbf{X}\mathbf{W}^{Q}_{h}$ is the query, $\mathbf{K}_{h}=\mathbf{X}\mathbf{W}^{K}_{h}$ is the key, $\mathbf{V}_{h}=\mathbf{X}\mathbf{W}^{V}_{h}$ is the value, $\mathbf{W}^{*}_{h}\in\mathbb{R}^{C\times d_{h}}$ ($*\in\{Q,K,V\}$), $\mathbf{W}^{O}\in\mathbb{R}^{C\times C}$, $H$ is the number of heads, $d_{h}=C/H$ is the feature dimension of head $h$, and the superscript is matrix transpose. Finally, FFN in Eq. (1) consists of two linear layers with a GELU [61] non-linearity in between, i.e.,
@@
Subsequently, we utilize $M_{i}-1$ SBT blocks to operate on the summarized bottleneck tokens to achieve efficient global spatiotemporal interactions. As shown in Fig. 2, an SBT block mainly consists of a Multi-Head Cross-Attention (MHCA) layer [62], an MHSA layer, and an FFN layer. Formally, it can be defined as follows:
@@
where the superscript $j\in\{1,...,M_{i}-1\}$ is the block index, $\mathbf{B}^{(0)}_{i}=\mathbf{B}_{i}$. The MHCA layer iteratively distills necessary information from the original input back into the bottleneck embeddings to avoid key information being lost during spatial compression. It is a variant of MHSA, whose query is the bottleneck embeddings while the key and value come from the original input to allow information flow from the latter to the former, i.e.,
@@
Compared with a standard Transformer block, the computational complexity of an SBT block is reduced from the quadratic $O(S^{2}T_{i}^{2})$ to $O(G(G+S)T_{i}^{2})$ thanks to the introduction of spatial bottleneck tokens. Considering that $G$ is a small constant, an SBT block thus enjoys linear computational complexity with respect to the spatial size $S$. Note that, empirically, we have $G\approx S$ during pre-training and $G\ll S$ during fine-tuning (Sec. 4.1). Therefore, the computational costs of pre-training remain approximately unchanged (actually, slightly decrease) and large computations can be reduced during downstream fine-tuning, allowing both efficient pre-training and fine-tuning (Sec. 4.2.2).